Graham, a physics enthusiast, is asked by his instructor to visit a nearby carpentry. He observed
there a hand–held powered reciprocating saw, say a sabre saw, which performs SHM in a vertical plane with
an angular frequency of 350 rad s−1
. Assume that the tip of the toothed blade of the saw is at x = 0 with
positive velocity when t = 0 and the amplitude of SHM is 2.40 cm. Write an equation, as a function of time,
which shows the position of the tip of the toothed blade. Also calculate that how much will the blade take to
travel from x = 0 to x = 1.20 cm? To 2.40 cm?
1
Expert's answer
2019-10-29T10:55:58-0400

equation of SHM is given by:


"x=a\sin (\omega t)"

Here,


"a=2.4\ cm\\ \omega=350\ red\ sec^{-1}\\So,\\x=2.4\sin(350t)"

(a) Time from x=0 to x=1.2 cm -


"1.2=2.4\sin(350t)\\or\\\sin(350t)=\frac{1}{2}\\or\\350t=\frac{\pi}{6}\\or\\t=0.0015\ sec"

(b) Time from x=0 to x = 2.4 cm


"2.4=2.4\sin(350t)\\or\\\sin(350t)=1\\or\\350t=\frac{\pi}{2}\\or\\t=0.0045\ sec"